Mary is presently twice as old as Maggy. Ten years ago, Maggy's age was one-third of Mary's then. Ten years from now, Maggy's age will be three-fifths of Mary's age then. How old is each?
Answer by stanbon(48545)

Mary is presently twice as old as Maggy. Ten years ago, Maggy's age was one-third of Mary's then. Ten years from now, Maggy's age will be three-fifths of Mary's age then. How old is each?
---------------------
NOW DATA:
Let Maggy's age be x yrs old; Mary's age is 2x yrs old
--------------------------------
Ten Yrs Ago DATA:
Maggy's age then = x-10 yrs old; Mary's age then 2x-10 yrs old
---------------------------
EQUATION:
x-10 = (1/3)(2x-10)
x-10 = (2/3)x-(1/10)
(1/3)x = 9.9
x = 29.7 yrs old (Maggy's age NOW)
2x = 59.4 yrs old (Mary's age NOW)
============
Cheers,
Stan H.